A typical atom has a diameter of about 4.56 x10 -10 m
a) What is the diameter in Inches and µm?
b) Approximately how many atoms are there along a 1.0 cm line?
1
Expert's answer
2020-05-07T10:22:45-0400

a) 1 inch = 2.54 cm, therefore "1 m = \dfrac{100\,cm}{2.54 cm/inch}." Next,

"4.56 \times 10^{-10} m = 4.56 \times 10^{-10} \times \dfrac{100\,cm}{2.54 cm/inch} = 1.80\times 10^{-8}\, inch."

1 "\mu m = 10^{-6}m," so "\;\; 4.56 \times 10^{-10} m = 4.56\times 10^{-4} \times 10^{-6}\, m = 4.56\times 10^{-4} \,\mu m."

b) To calculate the number of atoms we should divide 1.0 cm by the diameter of a single atom, so we get

"N=\dfrac{1.0\,cm}{4.56\times10^{-10}\,m} = \dfrac{0.010\,m}{4.56\times10^{-10}\,m} = 21929824 \approx 2.2\cdot10^7."

The number is approximately 22 mln.